OpenStudy (anonymous):

Find the arc length of the curve: y=(1/2)((e^x)+(e^-x))

OpenStudy (anonymous):

I know that (1/2)((e^x)+(e^-x)) = coshx but I am not sure what to do after that.

OpenStudy (anonymous):

if i recall correctly the arc length is given by \[\int_a^b\sqrt{1+(f'(x)})^2dx\]

OpenStudy (anonymous):

Could you help me with the derivation? d/dx of e^x is just e^x, right? So would e^-x be e^-x?

OpenStudy (anonymous):

no it would be \(-e^{-x}\) by the chain rule

OpenStudy (anonymous):

so your derivative would be \[\frac{e^x-e^{-x}}{2}\]

OpenStudy (anonymous):

or with hyperbolic function, you have cosh and the derivative is sinh

OpenStudy (anonymous):

and by coincidence, \(1+\sinh^2(x)=\cosh^2(x)\) i believe

OpenStudy (anonymous):

Yes, got the same thing. The hyperbolic method seems like it would be easier, however I am not very well versed in them.

OpenStudy (anonymous):

i think that is the gimmick here, because these problems are always cooked up to give you something you can actually integrate

OpenStudy (anonymous):

so you take the derivative of cosh, get sinh, then square and add one and you get \(\cosh^2(x)\)

OpenStudy (anonymous):

take the square root and you get \(\cosh(x)\) again

OpenStudy (anonymous):

i assume you have limits of integration too right?

OpenStudy (anonymous):

Yes, the arc is from x=0, and x=1.